The Amount of an Endowment

A university alumni group wishes to provide an annual scholarship in the amount of $1400 beginning next year. If the scholarship fund will earn an interest rate of 4%/year compounded continuously, find the amount of the endowment the alumni are required to make now. (Assume that the beginning of next year corresponds to

t = 0.)
Answer #1
EAR =[ e^(Annual percentage rate) -1]*100
Effective Annual Rate=(e^(4/100)-1)*100
Effective Annual Rate% = 4.0811
Future value = present value*(1+ rate)^time
1400 = Present value*(1+0.040811)^1
Present value = 1345.1
Using Calculator: press buttons "2ND"+"FV" then assign
FV =-1400
I/Y =4.0811
N =1
PMT = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(0.040811,1,,1400,)